User name Markowitz
Log entry time 23:58:15 on December21,2003
Entry number 114975
runs 1504-1515 all RICH TH delay scan (see Bodo's HALOG)
18:30 change to H2O for waterfall calibration. Magnets set to 2.030 GeV momentum (cycled Q2 and Q3).
Run 1516, pump speed 3000,
run 1517, pump speed 3500,
run 1518 junk
run 1519, pump speed 4000,
run 1520, pump speed 2500, RICH at 2120 V
Were not able to keep up to DAQ on analysis of H2O calibration.
Run 1521 same pump speed, RICH at 2150 V, prescales have changed
Run 1522 junk
run 1523 same pump speed, RICH at 2180 V.
run 1524 same pump speed, RICH at 2090 V.
run 1525 same pump speed, RICH at 2060 V.
run 1526 same pump speed, RICH at 2000 V.
run 1527 same pump speed, RICH at 1900 V. junk, pedestals not loaded
run 1528 same pump speed, RICH at 1900 V.
run 1529 same pump speed, RICH at 2210 V.
run 1530 thick carbon, prescale factors and conditions from 1516ff
set HRS both to 1.010 GeV
Doing EP measurement (target appears to be working, see Bodo's HALOG)